OpenStudy (anonymous):

can someone explain me how to do this?? Celine has a bottle that contains 20% of milk and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of milk and water in the bottle. Define the variable used in the equation. (5 points) Part B: How many liters of milk are present in the bottle? Show your work. (5 points)

OpenStudy (anonymous):

@SolomonZelman ?

OpenStudy (anonymous):

@Koikkara @KlOwNlOvE @hba @ObeySoph▲

OpenStudy (anonymous):

well what do you know about how many liters of milk is in it

OpenStudy (anonymous):

20%

OpenStudy (anonymous):

alright so we are going to set up proportions do you know about those?

OpenStudy (anonymous):

isn't it like this/this=this/this??? just a rough example im not so sure

OpenStudy (anonymous):

sorry internet is cutting in and out but yes

OpenStudy (anonymous):

so percentage/liter is how we will set up

OpenStudy (anonymous):

\[\frac{ 20 }{ 1 } \frac{ 80 }{ x }\]

OpenStudy (anonymous):

but it says equation and not proportion

OpenStudy (anonymous):

do you understand where i got the 80 from?

OpenStudy (anonymous):

a proportion is a way to set up equations

OpenStudy (anonymous):

no I don't get where the 80 is coming from

OpenStudy (anonymous):

alright in a bottle theres 100% liquid 20% is milk so 80% is water

OpenStudy (anonymous):

oh yeah that makes sense

OpenStudy (anonymous):

ok can you cross multiply that?

OpenStudy (anonymous):

80*1= 20*x=

OpenStudy (anonymous):

4

OpenStudy (anonymous):

yes x=4

OpenStudy (anonymous):

and that would be a proportion equation

OpenStudy (anonymous):

but that doesn't really make sense to what I have read about this specific question. hold up ill send you the link of how others explained it. I don't get how they explain it though

OpenStudy (anonymous):

oh yes theyre explaining the amount of milk in the liter i gave percentage for water

OpenStudy (anonymous):

im not sure then not understanding how they explained

OpenStudy (anonymous):

oh ok but thank you for trying

OpenStudy (anonymous):

do you know someone who could help me??

OpenStudy (anonymous):

if you ask ganeshi they might help im not seeing them online tho

OpenStudy (anonymous):

oh ok ill just tag them and hope they'll reply

OpenStudy (anonymous):

@ganeshi

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

@Whitemonsterbunny17

OpenStudy (anonymous):

Heelp???

OpenStudy (anonymous):

@uri @Olaelyamany @LazyBoy @superhelp101

OpenStudy (jdoe0001):

well, seems to me the amount of water in the bottle is given "The bottle has 1 liter of water."

OpenStudy (anonymous):

yeah I know that the whole question though confuses me. idk why

OpenStudy (jdoe0001):

keep in mind that Celine has a bottle, 20% of that is milk, the rest, 80%, is water the bottle has a total of 1liter of water, meaning that 80% of the bottle is 1liter because 80% of the bottle, the rest, is water

OpenStudy (anonymous):

ok but idk how I would set up the equation. it is a little confusing. usually im good at that but this is just idk how to say it without being mean about it c; lol

OpenStudy (jdoe0001):

ok... so we know that 20% is milk and that 80% is water, and that 80% is 1liter so \(\large \begin{array}{ccllll} liters&\% \\\hline\\ 1&80\\ x&20 \end{array} \bf \implies \cfrac{1}{x}=\cfrac{80}{20}\)

OpenStudy (jdoe0001):

and if you solve for "x", that'd be how much the 20% is in liters

OpenStudy (anonymous):

ok thank you so much

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

right idea but set it up wrong sorry >.<

OpenStudy (anonymous):

its alright